OpenStudy (anonymous):

between which consecutive integers does the square root 90 lie?

OpenStudy (joannablackwelder):

1^2=1 2^2=4, 3^2=9, 4^2=16, 5^2=25, 6^2=36, 7^2=49, 8^2=64, 9^2=81, 10^2=100

OpenStudy (joannablackwelder):

So, that means the square root of 81=9

OpenStudy (joannablackwelder):

And the square root of 100 is 10.

OpenStudy (joannablackwelder):

The square root of 90 would be somewhere in between 9 and 10.

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (anonymous):

thanks! :)